OpenStudy (anonymous):

I need help answering this question. C: What is the quadratic function that is created with roots at 2 and 4 and a vertex at (3, 1)?

OpenStudy (anonymous):

f(x) = -3x2 + 12x?

OpenStudy (anonymous):

it may be f(x)= 1 - (x-3)^2

OpenStudy (anonymous):

It can't be because my choices are. f(x) = -3x2 + 12x f(x) = 3x2 − 12x f(x) = 2x2 − 8x f(x) = -2x2 + 8x

OpenStudy (anonymous):

f(x) vanishes at x=2 and x=4 and has the vertex at 3,1

OpenStudy (anonymous):

if so it may be f(x) = 3x2 − 12x

OpenStudy (anonymous):

Wait no! Omg! Fail on my part! f(x) = -x2 + 6x − 8 f(x) = x2 − 6x − 8 (x) = -x2 + 6x − 8 (x) = x2 − 6x + 8

OpenStudy (anonymous):

Those are the REAL choices

OpenStudy (anonymous):

i think it is the first one the equations is given by f(x) = a(x-2)(x-4) to determine (a) we must use the vertex (3,1)